2) White light is being analyzed by a diffraction grating (100 lines/mm). In the interference pattern, the third order blue (wavelength=400 nm) appears 10 cm from the central maxima. How far away is the screen placed from the grating?

a) 0.83 m b) 6.52m c) 0.21m d) 8.33m

Do you agree or disagree with the statement: "Single-slit interference shouldn't be possible if light was wave; for waves to interfere- you need at least two slits"

a) agree b)disagree
Answer #1

a)
Spacing between diffraction grating, d = 1 mm/100
= 10 x 10-6 m

For the third order diffraction maxima,
3 x = d sin
sin = 3/d
= [3 x (400 x 10-9)] / (10 x 10-6)
= 0.12

= sin-1(0.12)
= 6.89 degrees

tan = x/D
Where x is the distance from the central maxima to the third maxima and D is the distance of the screen from the grating.
D = x/tan
= 10/tan(6.89)
= 0.83

b)
Disagree.
For the interference to occur, single slit is enough. Light comes from the right edge of the single slit may have travelled a different distance compared with the light comes from the left edge and diffraction will happen